# RedTeamCUA: Realistic Adversarial Testing of Computer-Use Agents in Hybrid Web-OS Environments

### v4

@@ Abstract
Computer-use agents (CUAs) promise to automate complex tasks across operating systems (OS) and the web, but remain vulnerable to indirect prompt injection, where attackers embed malicious content into the environment to hijack agent behavior. Current evaluations of this threat either lack support for adversarial testing in realistic but controlled environments or ignore hybrid web-OS attack scenarios involving both interfaces. To address this, we propose RedTeamCUA, an adversarial testing framework featuring a novel hybrid sandbox that integrates a VM-based OS environment with Docker-based web platforms. Our sandbox supports key features tailored for red teaming, such as flexible adversarial scenario configuration, and a setting that decouples adversarial evaluation from navigational limitations of CUAs by initializing tests directly at the point of an adversarial injection. Using RedTeamCUA, we develop RTC-Bench, a comprehensive benchmark with 864 examples that investigate realistic, hybrid web-OS attack scenarios and fundamental security vulnerabilities. Benchmarking current frontier CUAs identifies significant vulnerabilities: Claude 3.7 Sonnet | CUA demonstrates an Attack Success Rate (ASR) of 42.9%, while Operator, the most secure CUA evaluated, still exhibits an ASR of 7.6%. Notably, CUAs often attempt to execute adversarial tasks with an Attempt Rate as high as 92.5%, although failing to complete them due to capability limitations. Nevertheless, we observe concerning high ASRs in realistic end-to-end settings, with the recent Claude 4.5 Opus | CUA and strongest-to-date Claude 4.6 Opus | CUA exhibiting a concerning ASR of 83% and 50%, respectively. This indicates that CUA threats can already result in tangible risks to users and computer systems. Overall, RedTeamCUA provides an essential framework for advancing realistic, controlled, and systematic analysis of CUA vulnerabilities, highlighting the urgent need for robust defenses to indirect prompt injection prior to real-world deployment.
[Uncaptioned image] Website https://osu-nlp-group.github.io/RedTeamCUA
[Uncaptioned image] Code https://github.com/OSU-NLP-Group/RedTeamCUA
@@ 2.2 A Critical Need for a Hybrid Environment Sandbox
Given these limitations, we highlight the critical need for a hybrid sandbox that enables realistic and interactive adversarial evaluation across secure web and OS environments (§ 3) and a large-scale adversarial benchmark with broad coverage of severe adversarial scenarios (§ 4). In addition, we provide a detailed comparison of existing work in Table 6 of Appendix C.
@@ 4.1 Benign Goal Formulation
To align with our focused CUA use cases (described in §2.1) where users fetch online knowledge for local execution, we define benign goals across three categories: (1) Software Installation, where the agent installs tools, libraries, or packages found online, (2) System Configuration, where the agent configures or customize local system settings, and (3) Project Setup, where the agent downloads a codebase or dataset aligned with the user’s goals. We create 3 distinct benign goals per category using web environments in RedTeamCUA, resulting in 9 total goals (Appendix A.2). Beyond this, to simulate varying levels of user expertise occurring in real scenarios, we design two instantiations of benign instructions: General, where the user provides vague, high-level instructions, and Specific, where the user provides more detailed instructions based on their domain knowledge.
@@ 4.2 Adversarial Attack Formulation
Threat Model: Our threat model focuses on indirect prompt injection, where malicious content is embedded in web environments to manipulate an agent to deviate from its benign goal and perform a harmful task. We assume realistic attacker constraints: the attacker cannot access or modify the user’s original instruction, the agent’s prompts, components or model weights, and can only inject content into locations on a webpage where textual input is typically permitted (e.g., Forum comments, RocketChat messages, shared OwnCloud files). Unlike prior work that assumes attackers have unrealistically full webpage or OS access (Zhang et al., 2024; Liao et al., 2025; Chen et al., 2025a; Boisvert et al., 2025), our threat model reflects the real-world scenario in which platforms have strict UI design standards and access controls preventing unauthorized web modifications. Due to this, we focus on realistic, text-based injection within editable web content rather than misleading visual pop-ups or arbitrary UI manipulation from attackers. Due to the attacker’s lack of knowledge of the user’s instruction, we assume an adversarial strategy where the attacker blends their injection into the environment context to match anticipated user queries for a given web page. For example, the attacker may target CUAs on a Forum page related to software installation using an adversarial comment that couples harmful instructions with legitimate installation steps (shown in Figure 1). We explore scenarios where a user’s benign task aligns with the attacker’s contextualized injection, allowing us to assess attack viability under high-risk adversarial conditions (examples in Appendix A.4).
@@ 5.1 Setup
$\bullet$ Adapted LLM-based Agents include powerful LLMs adapted for computer use through generic agentic scaffolding. For this category, we evaluate GPT-4o (Hurst et al., 2024), the base versions of Claude 3.5 Sonnet (v2) (Anthropic., 2024c) and Claude 3.7 Sonnet (Anthropic., 2025a). For these agents, we follow the default agentic scaffolding provided by OSWorld, which uses pyautogui for Python-based execution of mouse and keyboard commands and provides necessary contextual information within the system prompt (see Appendix I).
$\bullet$ Specialized Computer-Use Agents are designed specifically for computer use, featuring training for GUI perception (OpenAI., 2025b; Anthropic., 2024b) and reasoning (OpenAI., 2025b) and incorporating tailored computer-use tools (Anthropic., 2024a). For this category, we mainly evaluate Operator (OpenAI., 2025b) and the computer-use versions of Claude 3.5 Sonnet (v2) and Claude 3.7 Sonnet (Anthropic., 2024a). Since their native action formats are often incompatible with the pyautogui-based execution in OSWorld, we employ GPT-4o as an auxiliary LLM to convert their native action outputs into executable pyautogui commands (prompt shown in Appendix I).
@@
Evaluation Metrics: To evaluate the success of both benign and adversarial tasks, we adopt example-specific execution-based evaluators to compute Success Rate (SR) and Attack Success Rate (ASR), respectively. This contrasts with OS-Harm’s (Kuntz et al., 2025) reliance on an automated LM judge which can itself be misled by prompt injections, undermining its reliability for adversarial testing. Evaluation based on executable scripts helps ensure robustness against different agent trajectories leading to the same outcome (Xie et al., 2024; 2025). Nonetheless, execution-based evaluation alone may fail to capture an agent’s susceptibility to indirect prompt injection, as an agent might be successfully misled to attempt an adversarial goal, and only fail to fully complete it due to limited capabilities. To address this, we introduce Attempt Rate (AR), a fine-grained LLM-as-a-Judge metric (GPT-4o, prompt in Appendix I) to assess whether an agent attempts to pursue an adversarial goal within the trajectory, regardless of harmful task completion. Together, ASR and AR balance reliable evaluation with broader coverage of harmful behaviors.
@@ C I A C I A C I A
Additional Details: We conduct sanity checks to ensure that all evaluated CUAs can successfully finish the benign tasks without injection prior to experimentation. Unless otherwise specified, agents operate with screenshot observations to align with how humans navigate computer-use environments (Gou et al., 2025; Qin et al., 2025) and are evaluated under the Decoupled Eval setting, where pre-processed actions place the agent directly at the state containing adversarial injection to isolate robustness from navigational ability. Each example is tested three times, and an attack is counted as successful if any run succeeds. More details are in Appendix D.
5.2 Decoupled Eval Results
@@
Integrity goals involve simple actions (e.g., file deletion via sudo rm -rf /etc/security); Confidentiality goals demand complex multi-step operations (e.g, extracting, creating, and exfiltrating file content) which may depend on agent capabilities; and Availability goals present a range, from simple service disruption (e.g., sudo systemctl stop sshd) to more intricate resource exhaustion tasks (e.g., creating 10,000 1MB files to consume OS storage). Despite this, AR for Confidentiality goals remains high across platforms, indicating that future, more capable agents may make such dangerous privacy attacks feasible if manipulation risks are not mitigated. We also find that the same adversarial goal yields different ASRs across our available web platforms, with RocketChat consistently resulting in the highest ASR compared to the other two platforms. We attribute this to two factors: (1) The perceived trustworthiness of the content source in a given environment could impact attack success, as agents might implicitly place higher trust in direct user messages compared to less trusted content like public forum comments. (2) The naturalness and plausibility of performing the harmful action within the context of the specific web platform could impact attack susceptibility. This is evidenced by the notably higher ASR and AR for Confidentiality goals within RocketChat, as sending data is more aligned with the platform’s inherent usage as a messaging tool and is easier to perform compared to our other web environments. These findings suggest that both web platform characteristics and adversarial goal types jointly influence attack success, leaving room for further exploration in future CUA red-teaming and defense work.
Adapted LLM-based Agents vs. Specialized Computer-Use Agents: Our results in Table 1 and Figure 10 (in Appendix H.5) reveal an interesting contrast between Specialized Computer-Use Agents developed by OpenAI and Anthropic. While Adapted LLM-based Agents from both organizations demonstrate relatively high ASR and AR, Operator achieves a substantial reduction ($58.62\%$ ASR) in these metrics to stand out as the most secure CUA, while the CUA version of Claude 3.7 Sonnet unexpectedly shows increased susceptibility ($3.66\%$ ASR) compared to its base counterparts. Operator specifically benefits from its built-in confirmation and safety check mechanisms, indicating that future CUA defenses can benefit from features introducing explicit user permission before executing critical or high-risk actions. However, the average ASR and AR remain high for the Operator (w/o checks) variant in the absence of reliable human supervision at $\sim$31% and $\sim$48% respectively (shown in Table 1). This exposes an inherent trade-off between agent autonomy and security: while human oversight can provide critical guardrails, truly autonomous CUAs require more robust internal safety mechanisms to independently detect and refuse harmful instructions.
@@ 6 Analysis
Defenses: Following Chen et al. (2025d), we categorize existing prompt injection defenses into two types: (1) System-Level Defenses, such as the recent LlamaFirewall (Chennabasappa et al., 2025) and PromptArmor (Shi et al., 2025), which prevent exposure to injection through additional detectors, monitors, or prompting; and (2) Model-Level Defenses, such as the open-source secure foundation model Meta SecAlign (Chen et al., 2025d), which build in security by training a foundation model to prioritize trusted instruction over untrusted data. While our evaluated proprietary CUAs already incorporate both levels of defenses according to OpenAI. (2025a) and Anthropic. (2025a), we further assess four defense methods from both categories for more comprehensive evaluation, using 50 high-risk examples from RTC-Bench that achieved the highest ASR across agents in §5 (Appendix G). For System-Level Defenses, we find that (1) LlamaFirewall and PromptArmor perform poorly in our setting, with the best variant detecting only 30% of injections (Appendix G.1) and (2) a defensive system prompt that instructs the agent to detect injections and stick to the user’s original instruction provides insufficient protection, as ASR for Claude 3.7 Sonnet | CUA and Operator (w/o checks) remain near 50% (Appendix G.2). For Model-Level Defenses, Meta SecAlign still follows malicious instructions in about half of the tasks (Appendix G.3). These findings underscore the need for more effective defenses specifically tailored to CUAs to ensure safe and secure real-world deployment.
Additional Analysis: We perform additional ablations (Appendix H), including the following results: (1) ASR can be reduced with more specific benign instructions and CUA use cases requiring less autonomy (H.1). (2) The success of different adversarial injection modalities (Code vs. Language) can be directly affected by web platform characteristics (H.1). (3) Observations using accessibility (a11y) trees can reduce ASR but may hurt benign task SR, suggesting a capability-safety trade-off (H.2). (4) Additional injection doesn’t impair the CUAs’ utility as demonstrated by the identical SR and ASR under the End2End setting (H.4). (5) A certain amount of adversarial risks consistently persist across all three runs, necessitating deeper exploration into preventing them(H.5).
@@ 7 End2End Evaluation
While Decoupled Eval isolates adversarial robustness from capability limits, it creates a gap with real-world end-to-end use of CUAs. To bridge this gap, we evaluate the same 50 tasks used in the previous defense experiment in the End2End setting, where CUAs start from the initial task state rather than the page containing adversarial injection. Additionally, beyond the CUAs evaluated in §5, we also incorporate the recently released CUAs with more advanced capabilities in computer-use scenarios to assess and track their robustness against prompt injection risks. Concretely, we include Claude 4 Opus | CUA (Anthropic., 2025b), Claude 4.5 Sonnet | CUA (Anthropic., 2025c), Claude 4.5 Opus | CUA (Anthropic., 2025e) and Claude 4.6 Opus | CUA (Anthropic., 2026) which achieves the SOTA performance on OSWorld.
@@
As shown in Figure 3, both Operator (w/o checks) and Claude 3.7 Sonnet | CUA demonstrate notable ASR in End2End evaluation, reaching an end-to-end ASR of 42% and 50%, respectively. However, our manual inspection of CUA outcomes for each adversarial task (Table. 3) reveals that both CUAs frequently fail to navigate to the site of injection or fully complete the injected task despite attempting, suggesting that risks could increase further with improved capabilities. Most alarmingly, Claude 4.5 Sonnet | CUA exhibits an increased ASR of 60% and Claude 4.5 Opus | CUA exhibits the highest ASR of 83% among all evaluated CUAs. Our manual inspection reveals that their enhanced capabilities enable them to overcome the navigational and task completion limitations observed in weaker CUAs, amplifying the success of adversarial attacks. Despite documented safeguards against prompt injection in these agents (Anthropic., 2025d; f), including reinforcement learning techniques and detection mechanisms, existing CUAs still remain alarmingly susceptible to adversarial manipulation. This reveals a concerning phenomenon: CUA capability improvements without sufficiently robust defenses could result in increased vulnerabilities and more serious risks.
Interestingly, we observe that the latest and strongest-to-date Claude 4.6 Opus | CUA shows reduced ASR compared to its predecessor, Claude 4.5 Opus | CUA. According to Table 3, in 25 out of 50 examples, Opus 4.6 correctly identifies malicious injections during its reasoning process, explicitly flagging them as adversarial attempts and subsequently adhering only to legitimate task instructions. In contrast, Opus 4.5 merely triggers such detection in only 3 cases. This represents a meaningful advancement in defending against adversarial risks. Nevertheless, Opus 4.6 still exhibits a 50% ASR, indicating that a substantial gap remains before reliable deployment can be achieved.
These findings together highlight the urgent need to further strengthen the adversarial robustness of CUAs to mitigate this significant vulnerability in future CUA releases alongside capability improvements.
@@
the Table 2. Such difference can largely be attributed to limitations in agent capabilities, e.g., the agent fails to navigate to the expected page. Nevertheless, attack success occurring within realistic, end-to-end evaluation highlights that real-world threats are no longer hypothetical and will only be amplified by CUA capability improvements in the near future. This underscores the value of our Decoupled Eval setting, allowing model developers to identify and proactively mitigate potential vulnerabilities before they manifest with more capable CUAs.
Table 3: A detailed breakdown of adversarial test outcomes for evaluation under the End2End setting based on manual inspection. Definitions for each possible outcome are provided in Appendix. H.6.

### v5

# arXiv:2505.21936v5[cs.CL] 01 Mar 2026

@@ Abstract
Computer-use agents (CUAs) promise to automate complex tasks across operating systems (OS) and the web, but remain vulnerable to indirect prompt injection, where attackers embed malicious content into the environment to hijack agent behavior. Current evaluations of this threat either lack support for adversarial testing in realistic but controlled environments or ignore hybrid web-OS attack scenarios involving both interfaces. To address this, we propose RedTeamCUA, an adversarial testing framework featuring a novel hybrid sandbox that integrates a VM-based OS environment with Docker-based web platforms. Our sandbox supports key features tailored for red teaming, such as flexible adversarial scenario configuration, and a setting that decouples adversarial evaluation from navigational limitations of CUAs by initializing tests directly at the point of an adversarial injection. Using RedTeamCUA, we develop RTC-Bench, a comprehensive benchmark with 864 examples that investigate realistic, hybrid web-OS attack scenarios and fundamental security vulnerabilities. Benchmarking current frontier CUAs identifies significant vulnerabilities: Claude 3.7 Sonnet | CUA demonstrates an Attack Success Rate (ASR) of 42.9%, while Operator, the most secure CUA evaluated, still exhibits an ASR of 7.6%. Notably, CUAs often attempt to execute adversarial tasks with an Attempt Rate as high as 92.5%, although failing to complete them due to capability limitations. Nevertheless, we observe concerning high ASRs in realistic end-to-end settings, with the recent Claude 4.5 Opus | CUA and strongest-to-date Claude 4.6 Opus | CUA exhibiting a concerning ASR of 83% and 50%, respectively. This indicates that CUA threats can already result in tangible risks to users and computer systems. Code and webpage are at https://osu-nlp-group.github.io/RedTeamCUA
@@ 2.2 A Critical Need for a Hybrid Environment Sandbox
Given these limitations, we highlight the critical need for a hybrid sandbox that enables realistic and interactive adversarial evaluation across secure web and OS environments (§ 3) and a large-scale adversarial benchmark with broad coverage of severe adversarial scenarios (§ 4). In addition, we provide a detailed comparison of existing work in Table 6 of Appendix D.
@@ 4.1 Benign Goal Formulation
To align with our focused CUA use cases (described in §2.1) where users fetch online knowledge for local execution, we define benign goals across three categories: (1) Software Installation, where the agent installs tools, libraries, or packages found online, (2) System Configuration, where the agent configures or customize local system settings, and (3) Project Setup, where the agent downloads a codebase or dataset aligned with the user’s goals. We create 3 distinct benign goals per category using web environments in RedTeamCUA, resulting in 9 total goals (Appendix B.2). Beyond this, to simulate varying levels of user expertise occurring in real scenarios, we design two instantiations of benign instructions: General, where the user provides vague, high-level instructions, and Specific, where the user provides more detailed instructions based on their domain knowledge.
@@ 4.2 Adversarial Attack Formulation
Threat Model: Our threat model focuses on indirect prompt injection, where malicious content is embedded in web environments to manipulate an agent to deviate from its benign goal and perform a harmful task. We assume realistic attacker constraints: the attacker cannot access or modify the user’s original instruction, the agent’s prompts, components or model weights, and can only inject content into locations on a webpage where textual input is typically permitted (e.g., Forum comments, RocketChat messages, shared OwnCloud files). Unlike prior work that assumes attackers have unrealistically full webpage or OS access (Zhang et al., 2024; Liao et al., 2025; Chen et al., 2025a; Boisvert et al., 2025), our threat model reflects the real-world scenario in which platforms have strict UI design standards and access controls preventing unauthorized web modifications. Due to this, we focus on realistic, text-based injection within editable web content rather than misleading visual pop-ups or arbitrary UI manipulation from attackers. Due to the attacker’s lack of knowledge of the user’s instruction, we assume an adversarial strategy where the attacker blends their injection into the environment context to match anticipated user queries for a given web page. For example, the attacker may target CUAs on a Forum page related to software installation using an adversarial comment that couples harmful instructions with legitimate installation steps (shown in Figure 1). We explore scenarios where a user’s benign task aligns with the attacker’s contextualized injection, allowing us to assess attack viability under high-risk adversarial conditions (examples in Appendix B.4).
@@ 5.1 Setup
$\bullet$ Adapted LLM-based Agents include powerful LLMs adapted for computer use through generic agentic scaffolding. For this category, we evaluate GPT-4o (Hurst et al., 2024), the base versions of Claude 3.5 Sonnet (v2) (Anthropic., 2024c) and Claude 3.7 Sonnet (Anthropic., 2025a). For these agents, we follow the default agentic scaffolding provided by OSWorld, which uses pyautogui for Python-based execution of mouse and keyboard commands and provides necessary contextual information within the system prompt (see Appendix J).
$\bullet$ Specialized Computer-Use Agents are designed specifically for computer use, featuring training for GUI perception (OpenAI., 2025b; Anthropic., 2024b) and reasoning (OpenAI., 2025b) and incorporating tailored computer-use tools (Anthropic., 2024a). For this category, we mainly evaluate Operator (OpenAI., 2025b) and the computer-use versions of Claude 3.5 Sonnet (v2) and Claude 3.7 Sonnet (Anthropic., 2024a). Since their native action formats are often incompatible with the pyautogui-based execution in OSWorld, we employ GPT-4o as an auxiliary LLM to convert their native action outputs into executable pyautogui commands (prompt shown in Appendix J).
@@
Evaluation Metrics: To evaluate the success of both benign and adversarial tasks, we adopt example-specific execution-based evaluators to compute Success Rate (SR) and Attack Success Rate (ASR), respectively. This contrasts with OS-Harm’s (Kuntz et al., 2025) reliance on an automated LM judge which can itself be misled by prompt injections, undermining its reliability for adversarial testing. Evaluation based on executable scripts helps ensure robustness against different agent trajectories leading to the same outcome (Xie et al., 2024; 2025). Nonetheless, execution-based evaluation alone may fail to capture an agent’s susceptibility to indirect prompt injection, as an agent might be successfully misled to attempt an adversarial goal, and only fail to fully complete it due to limited capabilities. To address this, we introduce Attempt Rate (AR), a fine-grained LLM-as-a-Judge metric (GPT-4o, prompt in Appendix J) to assess whether an agent attempts to pursue an adversarial goal within the trajectory, regardless of harmful task completion. Together, ASR and AR balance reliable evaluation with broader coverage of harmful behaviors.
@@ C I A C I A C I A
Additional Details: We conduct sanity checks to ensure that all evaluated CUAs can successfully finish the benign tasks without injection prior to experimentation. Unless otherwise specified, agents operate with screenshot observations to align with how humans navigate computer-use environments (Gou et al., 2025; Qin et al., 2025) and are evaluated under the Decoupled Eval setting, where pre-processed actions place the agent directly at the state containing adversarial injection to isolate robustness from navigational ability. Each example is tested three times, and an attack is counted as successful if any run succeeds. More details are in Appendix E.
5.2 Results
@@
Integrity goals involve simple actions (e.g., file deletion via sudo rm -rf /etc/security); Confidentiality goals demand complex multi-step operations (e.g, extracting, creating, and exfiltrating file content) which may depend on agent capabilities; and Availability goals present a range, from simple service disruption (e.g., sudo systemctl stop sshd) to more intricate resource exhaustion tasks (e.g., creating 10,000 1MB files to consume OS storage). Despite this, AR for Confidentiality goals remains high across platforms, indicating that future, more capable agents may make such dangerous privacy attacks feasible if manipulation risks are not mitigated.
We also find that the same adversarial goal yields different ASRs across our available web platforms, with RocketChat consistently resulting in the highest ASR compared to the other two platforms. We attribute this to two factors: (1) The perceived trustworthiness of the content source in a given environment could impact attack success, as agents might implicitly place higher trust in direct user messages compared to less trusted content like public forum comments. (2) The naturalness and plausibility of performing the harmful action within the context of the specific web platform could impact attack susceptibility. This is evidenced by the notably higher ASR and AR for Confidentiality goals within RocketChat, as sending data is more aligned with the platform’s inherent usage as a messaging tool and is easier to perform compared to our other web environments. These findings suggest that both web platform characteristics and adversarial goal types jointly influence attack success, leaving room for further exploration in future CUA red-teaming and defense work.
Adapted LLM-based Agents vs. Specialized Computer-Use Agents: Our results in Table 1 and Figure 10 (in Appendix I.5) reveal an interesting contrast between Specialized Computer-Use Agents developed by OpenAI and Anthropic. While Adapted LLM-based Agents from both organizations demonstrate relatively high ASR and AR, Operator achieves a substantial reduction ($58.62\%$ ASR) in these metrics to stand out as the most secure CUA, while the CUA version of Claude 3.7 Sonnet unexpectedly shows increased susceptibility ($3.66\%$ ASR) compared to its base counterparts. Operator specifically benefits from its built-in confirmation and safety check mechanisms, indicating that future CUA defenses can benefit from features introducing explicit user permission before executing critical or high-risk actions. However, the average ASR and AR remain high for the Operator (w/o checks) variant in the absence of reliable human supervision at $\sim$31% and $\sim$48% respectively (shown in Table 1). This exposes an inherent trade-off between agent autonomy and security: while human oversight can provide critical guardrails, truly autonomous CUAs require more robust internal safety mechanisms to independently detect and refuse harmful instructions.
@@ 6 Analysis
Defenses: Following Chen et al. (2025d), we categorize existing prompt injection defenses into two types: (1) System-Level Defenses, such as the recent LlamaFirewall (Chennabasappa et al., 2025) and PromptArmor (Shi et al., 2025), which prevent exposure to injection through additional detectors, monitors, or prompting; and (2) Model-Level Defenses, such as the open-source secure foundation model Meta SecAlign (Chen et al., 2025d), which build in security by training a foundation model to prioritize trusted instruction over untrusted data. While our evaluated proprietary CUAs already incorporate both levels of defenses according to OpenAI. (2025a) and Anthropic. (2025a), we further assess four defense methods from both categories for more comprehensive evaluation, using 50 high-risk examples from RTC-Bench that achieved the highest ASR across agents in §5 (Appendix H). For System-Level Defenses, we find that (1) LlamaFirewall and PromptArmor perform poorly in our setting, with the best variant detecting only 30% of injections (Appendix H.1) and (2) a defensive system prompt that instructs the agent to detect injections and stick to the user’s original instruction provides insufficient protection, as ASR for Claude 3.7 Sonnet | CUA and Operator (w/o checks) remain near 50% (Appendix H.2). For Model-Level Defenses, Meta SecAlign still follows malicious instructions in about half of the tasks (Appendix H.3). These findings underscore the need for more effective defenses specifically tailored to CUAs to ensure safe and secure real-world deployment.
Additional Analysis: We perform additional ablations (Appendix I), including the following results: (1) ASR can be reduced with more specific benign instructions and CUA use cases requiring less autonomy (I.1). (2) The success of different adversarial injection modalities (Code vs. Language) can be directly affected by web platform characteristics (I.1). (3) Observations using accessibility (a11y) trees can reduce ASR but may hurt benign task SR, suggesting a capability-safety trade-off (I.2). (4) Additional injection doesn’t impair the CUAs’ utility as demonstrated by the identical SR and ASR under the End2End setting (I.4). (5) A certain amount of adversarial risks consistently persist across all three runs, necessitating deeper exploration into preventing them(I.5).
@@ 7 End2End Evaluation
While Decoupled Eval isolates adversarial robustness from capability limits, it creates a gap with real-world end-to-end use of CUAs. To bridge this gap, we evaluate the same 50 tasks used in the previous defense experiment in the End2End setting, where CUAs start from the initial task state rather than the page containing adversarial injection. Additionally, beyond the CUAs evaluated in
@@
§5, we also incorporate the recently released CUAs with more advanced capabilities in computer-use scenarios to assess and track their robustness against prompt injection risks. Concretely, we include Claude 4 Opus | CUA (Anthropic., 2025b), Claude 4.5 Sonnet | CUA (Anthropic., 2025c), Claude 4.5 Opus | CUA (Anthropic., 2025e) and Claude 4.6 Opus | CUA (Anthropic., 2026) which achieves the SOTA performance on OSWorld.
As shown in Figure 3, both Operator (w/o checks) and Claude 3.7 Sonnet | CUA demonstrate notable ASR in End2End evaluation, reaching an end-to-end ASR of 42% and 50%, respectively. However, our manual inspection of CUA outcomes for each adversarial task (App. I.6) reveals that both CUAs frequently fail to navigate to the site of injection or fully complete the injected task despite attempting, suggesting that risks could increase further with improved capabilities. Most alarmingly, Claude 4.5 Sonnet | CUA exhibits an increased ASR of 60% and Claude 4.5 Opus | CUA exhibits the highest ASR of 83% among all evaluated CUAs. Our manual inspection reveals that their enhanced capabilities enable them to overcome the navigational and task completion limitations observed in weaker CUAs, amplifying the success of adversarial attacks. Despite documented safeguards against prompt injection in these agents (Anthropic., 2025d; f), including reinforcement learning techniques and detection mechanisms, existing CUAs still remain alarmingly susceptible to adversarial manipulation. This reveals a concerning phenomenon: CUA capability improvements without sufficiently robust defenses could result in increased vulnerabilities and more serious risks.
Interestingly, we observe that the latest and strongest-to-date Claude 4.6 Opus | CUA shows reduced ASR compared to its predecessor, Claude 4.5 Opus | CUA. According to App. I.6, in 25 out of 50 examples, Opus 4.6 correctly identifies malicious injections during its reasoning process, explicitly flagging them as adversarial attempts and subsequently adhering only to legitimate task instructions. In contrast, Opus 4.5 merely triggers such detection in only 3 cases. This represents a meaningful advancement in defending against adversarial risks. Nevertheless, Opus 4.6 still exhibits a 50% ASR, indicating that a substantial gap remains before reliable deployment can be achieved. These findings together highlight the urgent need to further strengthen the adversarial robustness of CUAs to mitigate this significant vulnerability in future CUA releases alongside capability improvements.
@@
Table 2. Such difference can largely be attributed to limitations in agent capabilities, e.g., the agent fails to navigate to the expected page. Nevertheless, attack success occurring within realistic, end-to-end evaluation highlights that real-world threats are no longer hypothetical and will only be amplified by CUA capability improvements in the near future. This underscores the value of our Decoupled Eval setting, allowing model developers to identify and proactively mitigate potential vulnerabilities before they manifest with more capable CUAs.
Table 3: A detailed breakdown of adversarial test outcomes for evaluation under the End2End setting based on manual inspection. Definitions for each possible outcome are provided in Appendix. I.6.